OpenStudy (james1769):

Rewrite f(x) = 2(x − 1)2 + 3 from vertex form to standard form.

OpenStudy (ltrout):

Standard form of that would be y=4x-1 I believe :)

OpenStudy (james1769):

f(x) =2x2 + 5 f(x) = 4x2− 8x + 7 f(x) = 2x2+ 4x + 5 f(x) = 2x2− 4x + 5

OpenStudy (james1769):

r the answers

OpenStudy (ltrout):

Okay one sec

OpenStudy (ltrout):

The final option is the answer. \[y=2x^{2}-4x+5\]

OpenStudy (james1769):

could u help with another?

OpenStudy (ltrout):

Sure! :)

OpenStudy (james1769):

Which of the following values "completes the square," or creates a perfect square trinomial, for x2 + 6x + ___? (2 points)

OpenStudy (ltrout):

Which math are you in?

OpenStudy (james1769):

algebra 2

OpenStudy (ltrout):

The constant term is missing in yours, so you would square half the coefficient of the linear term. So (6/2)^2 = 9 :)

OpenStudy (james1769):

thnks bro

OpenStudy (ltrout):

No problem! :)

OpenStudy (james1769):

got a coupl more

OpenStudy (james1769):

Using the completing-the-square method, rewrite f(x) = x2 + 4x − 1 in vertex form. (2 points)